OpenStudy (anonymous):

I'm sure this looks extremely easy.. But I cant figure it out. factoring 12x^2-10x-12=0

OpenStudy (anonymous):

Factor out a 2 to start.

OpenStudy (anonymous):

Okay.. I did that

OpenStudy (anonymous):

So then you divide by the 2 and end up with 6x^2 - 5x - 6 = 0

OpenStudy (anonymous):

Then you would use the AC method.

OpenStudy (anonymous):

AC method?

OpenStudy (anonymous):

is that where you take the coefficient of the first term and multiply it with the last term and find two multiples that add to equal the middle term?

OpenStudy (anonymous):

Yeah :D

OpenStudy (anonymous):

haha okay well I did that from the get go, but I'm left with 36?

OpenStudy (anonymous):

Sorry I was trying to fix a printer haha

OpenStudy (anonymous):

it's okay.. I totally understand. I'm just thankful someone replied!

OpenStudy (anonymous):

Actually I lied haha, wrong method sorry. Well, right method but in the wrong way. Step 1: Multiply your first term by your last term: 6 x 6 = 36

OpenStudy (anonymous):

Step 2: Find two integers when multiplied together get 36, but when added together equal your middle term: 5 -4 and 9

OpenStudy (anonymous):

OMG! I feel like the biggest idiot! I so had that I just didnt even thing about making the four negative wow. you probably think I'm so dumb lol

OpenStudy (anonymous):

Step 3: Rewrite the middle term by replacing the (5x) with the two numbers chosen. (-4 and 9) 6x^2 -4x +9x + 12 = 0

OpenStudy (anonymous):

Not at all :) Do you know where to go from there?

OpenStudy (anonymous):

Yes:] I do.. Thank you So much! And you better not go anywhere.. I'm sure I will need you again.. I have 50 more to go lol

OpenStudy (anonymous):

Haha alright :) I think I will write the rest of the steps anyway just in case anyone else needs to know how to do the AC method.

OpenStudy (anonymous):

I have another one..

OpenStudy (anonymous):

alright :)

OpenStudy (anonymous):

I'm sure you're going to think I'm dumb again.. lol

OpenStudy (anonymous):

I didn't think you were dumb in the first place..

OpenStudy (anonymous):

Did it not post?

OpenStudy (anonymous):

Nope haha

OpenStudy (anonymous):

2x^2+X-15=0 This seriously looks as easy as the first one.. Thats sweet of you to not think I'm dumb lol

OpenStudy (anonymous):

so what are you trying to find terms for?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Are you using the AC method? First term times last term? 2 X 15 So you are looking for terms that multiply to get 30 and add up to get your middle term, which in this case is 1

OpenStudy (anonymous):

It doesnt specify which method to use.. But I'm sure thats what it is lol

OpenStudy (anonymous):

Whenever you are trying to factor and you have a number in front of your x^2 term, you are going to want to use the AC method :)

OpenStudy (anonymous):

Ohtayyyy

OpenStudy (anonymous):

So do you know where to go from there?

OpenStudy (anonymous):

No lol

OpenStudy (anonymous):

oh.. haha so find numbers that when multiplied get 30 and add up to 1

OpenStudy (anonymous):

Sorry my laptop froze.. Anyways I got 6 and -5 ... Duh!

OpenStudy (anonymous):

Sorry, my internet went out, and my tutee came haha So, do you know where to go from there?

OpenStudy (anonymous):

Tutee? And yes I finished it.. I'm not going to do the rest of my paper though.. I'm just gonna turn it in as is haha

OpenStudy (anonymous):

Yeah, I am a math tutor. :) And you really should finish :/

OpenStudy (anonymous):

Aweee! That's awesome! You can tutor me too lol.. Anyways I'm excused from class because of a death in the family but I still have to turn it in before seven